Is the following sequence arithemtic? if,so idenitfy the common difference of -16,-22,-29,-37
Answer by Edwin McCravy(8999)

-16,-22,-29,-37
Let's find out if it's an arithmetic sequence:

2nd term - 1st term = (-22) - (-16) = -22+16 = -6

3rd term - 2nd term = (-29) - (-22) = -29+22 = -7

No, it's not because we didn't get the same thing
for both of those subtractions.

If it had been an arithmetic sequence those would 
have had to be the same common difference and also

4th term - 3rd term = (-37) - (-29) = -37+29 = -8

would have had to have been the same common
difference.  So this is not an arithmetic sequence.

Edwin